Suppose a firm, in discussions with customers, learns that customers identify eight factors that they evaluate for every order they receive from suppliers. The firm then finds that its performance is 86 percent on six of these factors and 83 percent on the other two factors, thus the firm has a perfect order performance of 27.9%. Suppose performance on two of the factors that had been 86 percent falls to 81 percent.

a. What is the firm’s probable perfect order performance? (Round your answer to 1 decimal place.)


b. By how many percentage points has perfect order performance decreased? Enter the absolute percentage point reduction, as opposed to percentage change. (Round your answer to 1 decimal place.)
Answer #1

(a)

Firm's probable perfect order performance = 0.86*0.86*0.86*0.86*0.81*0.81*0.83*0.83

Firm's probable perfect order performance = 0.24724 = 24.7%

(b)

Case 1, perfect order performance = 27.9%

Case 2, perfect order performance = 24.7%

Change in perfect order performance = (24.7 - 27.9) = (-3.2) = decrease of 3.2%

So, it is decreased by 3.2% .

PLEASE LEAVE A LIKE. IT REALLY HELPS ME A LOT. THANK YOU!!!